54007GD:

1). The price of a video game is $32.99 before tax. Andre bought the video game for 20% off. He then paid 8% sales tax on the discounted price. Part A: How much did Andre pay in sales tax? Show all work and steps in your solution. Part B: What is the total amount that Andre paid for the video game? Show all work and steps in your solution. 2). Jason left a bin outside in his garden to collect rainwater. He notices that 1 over 5 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

jhonyy9:

how many is 20°% of 32,99 ?

54007GD:

6.59800

54007GD:

so 32.9 minus 6.59800

54007GD:

@54007gd wrote:
so 32.9 minus 6.59800
32.99

54007GD:

which is 26.392

54007GD:

and 8% of that is 2.11136

54007GD:

wait dont we have to round the numbers first to the nearest cent?

jhonyy9:

this 8% of what you calculed there ?

Extrinix:

To find total price: \(32.99 - (0.20 \times 32.99) = xPrice\) \(xPrice - (0.08 \times xPrice) = tPrice\) To find the sales tax amount: \((0.08 \times xPrice) = sPrice\)